OpenStudy (anonymous):

Please help solve for: 1. log3(2x+3) – log3(x-2) = 2 2. log9(x-6) + log9(x+2) = 1

OpenStudy (anonymous):

wat grade r u in

OpenStudy (anonymous):

ok i can't help u but can u help me

OpenStudy (anonymous):

this use substitution to solve 6x-4y=54 -9x+2y=-69

OpenStudy (anonymous):

can u rephrase that

OpenStudy (anonymous):

Can someone please hlp me?